# MaFeRw: Query Rewriting with Multi-Aspect Feedbacks for Retrieval-Augmented Large Language Models

### v1

@@ Introduction
Retrieval-augmented generation (RAG) effectively addresses the issue of factual inaccuracy in large language models (LLMs) by integrating relevant retrieved information. It is widely used in various fields such as QA systems (Du and Ji 2022; Siriwardhana et al. 2023; Sharma et al. 2024), content generation (Zhao et al. 2024; Su et al. 2024), and virtual agents (Yang, Yue, and He 2023; Schick et al. 2023; Zhang 2023). In real-world RAG systems, users describe their information needs using spoken language and employ multi-turn dialogues for retrieval and generation. Therefore, current queries often contain spoken ellipses and ambiguous references from dialogue contexts (Mao et al. 2023b), making it challenging for RAG systems to accurately understand user intent (Ye et al. 2024). In light of this challenge, effective query rewriting is necessary to better describe user information needs and ensure the accuracy of retrieval and generation in RAG.
@@
Currently, RAG mainly performs context-based query rewriting to improve the quality of response generation, but the resulting enhancement is minimal due to the lengthy process from query rewriting to response generation. Although context-based query rewriting can directly affect the retrieval results, this impact is difficult to transmit to the final generated results through this loose-coupling retrieval and generation framework of RAG. As shown in Figure 1, the T5 rewriter (Lin et al. 2020) trained on context-based rewrite dataset fails to effectively capture the user’s intent. To address this issue, Ma et al. (2023) try to use reinforcement learning (RL) with generation feedback to assist the rewriter. They optimize the rewriter with policy gradient and use the accuracy and ’hit’ rate of RAG inference’s answers to form the reward. However, this reward is sparse and provides little guidance in most cases, resulting in unstable training and generation results of their model. Additionally, it is inefficient to directly use the full RAG inference to obtain RL feedback. Because RL typically requires a large number of training steps to achieve good results, and each inference in RAG takes a long time. Thus, if a stable and efficient method for training the rewriter can be designed, it will undoubtedly further enhance the effectiveness of RAG.
@@
In dialogue systems, user utterances often contain omissions and ambiguous references. Therefore, a rewriting model is needed to resolve these ambiguities in the current query and recover missing elements (e.g., anaphora) from the context to better articulate the user’s information needs. Current research on query rewriting mainly focuses on conversational search tasks, conversational QA tasks, and RAG tasks.
In conversational search tasks, researchers try to explore effective query rewriting methods to enhance the retrieval process. Some studies focus on extracting user intents from the dialogue context. For example, Lin et al. (2020) use human rewrites as labels to train a sequence-to-sequence model as the rewriter. Qian and Dou (2022) propose query rewriting explicitly highlights relevant terms in the query context, facilitating contextual modeling and thereby improving the quality of the learned contextual query embeddings. In (Ye et al. 2023), LLM is used as the rewriter and is prompted to complete a ”rewrite-then-edit” process based on the conversation history. Others propose to introduce the retrieval feedback to make the rewriting model better serve the retrieval task. Mo et al. (2023) and Mao et al. (2023a) both take the feature similarity between the query and the target passage into the optimization objective when training the rewriting model, and Wu et al. (2022) use the retrieval effect as a reward for reinforcement learning. The rewritten queries are encouraged to achieve better retrieval performance.
In conversational QA tasks, mainstream query rewriting studies (Kim et al. 2021; Chen et al. 2022) focus on feeding the generated results back into the rewriting model to improve the accuracy of the final answers. Kim et al. (2021) train the QA model by regularizing the consistency of generated answers, so that the answers predicted by the model according to the original questions were similar to the answers generated according to the rewritten questions. Chen et al. (2022) use the effect of generated answers as a reward for reinforcement learning, to improve the effect of rewriting on the final generated results.
In conversational RAG task, Ma et al. (2023) firstly create a pseudo-dataset for supervised warm-up training, then further optimize the rewriter using a policy gradient RL framework, using the accuracy and ’hit’ rate of LLMs’ answers to form the reward. However, for RAG tasks, focusing solely on the retrieval process results in rewrites unsuitable for prompting LLMs during generation. Conversely, considering only the feedback from the final generated results can lead to unstable training of the rewriting model due to the sparse reward signal. Therefore, we propose MaFeRw, which combines multi-aspect reward feedbacks to produce more stable and satisfactory responses.
@@
For MaFeRw, the rewriter is trained to transform the current query $q^{k}$ into $q_{R}^{k}$ utilizing the dialogue history $h^{k}$, which enables the RAG system to generate more satisfactory results. Firstly, we use manual rewriting data to train a T5 model as the rewriter’s initialization. Secondly, we design three metrics as RL feedback: the similarity between the rewritten query $q_{R}^{k}$ and the gold document $d_{+}^{k}$, the ranking metrics of similarity between retrieved documents $D^{k}$ and the ground truth $G^{k}$, and ROUGE between the generation $g^{k}$ and the ground truth $G^{k}$. Inspired by RLAIF, we train three kinds of reward models for the above metrics on three corresponding paired data sets, as shown in Figure 2(a). Considering the obvious improvement of manual rewriting in the generation, ROUGE between manual rewrites and model rewrites is used as another feedback. Finally, we combine scores of these reward models and the rewritten ROUGE to feedback PPO training, exploring the optimal query rewriting strategy, as shown in Figure 2(b).
@@
Since user needs are substantially embedded within the ground truth $G$, and the order of retrieved documents $D$ also affects the quality of generations, we hope that the rewritten query can assist the retrieval model in prioritizing the document that exhibits high semantic relevance to the ground truth. Therefore, we propose a ranking metric $m_{D}$ using $G$ and $D=\{d_{1},\cdots,d_{l}\}$. Specifically, we transform the ground truth $G$ into a dense vector $v_{G}$ using the retrieval embedding model. Then the CS value is calculated between $v_{G}$ and dense vectors of retrieved documents $\{v_{d_{i}}\}_{i=1}^{l}$. For each $d_{i}$, let $1/i$ be its rank score. The metric $m_{D}$ is computed by summing the product of each document’s rank score and its CS value with $v_{G}$, formulated as $m_{D}=\Sigma_{i=1}^{l}1/i\cdot\mathrm{CS}(v_{G},v_{d_{i}})$. A higher $m_{D}$ indicates that the $d_{i}$ with greater similarity to $G$ is ranked higher, thereby facilitating the generation of better responses.
@@
Due to the prolonged inference time required for each iteration of RAG and the substantial number of training steps often necessary for RL to achieve satisfactory results, directly employing the complete RAG process for RL can result in an extensive training time. Inspired by RLAIF, we train three types of reward models for the aforementioned metrics on three corresponding paired datasets. The reward model is based on the T5-base model with a value head, possessing significantly fewer parameters than the generation model Llama2-13b. During RL training, the reward is derived from scores of reward models rather than using the complete RAG process, thereby reducing RAG inference time.
@@
here ”+” denotes string concatenation.
When training the reward model, we adopt a method similar to Wang et al. (2024). By following the Bradley-Terry model (Bradley and Terry 1952), we formulate a preference distribution by employing the reward model $r_{\psi}$ as outlined below:
@@
where $\eta$ is a coefficient governing the magnitude of KL penalty. The initial policy model is the initialized rewriter.
@@
We compare MaFeRw with four conversational RAG methods: raw RAG (Lewis et al. 2021), T5-base rewriter (Lin et al. 2020), RL-base rewriter (Ma et al. 2023), and ConvGQR (Mo et al. 2023). Raw RAG is the baseline RAG model that does not perform any query rewriting. It concatenates the dialogue history and the current query to perform retrieval and response generation. T5-base rewriter employs manual rewrites to train a T5 model serving as the rewriter. It rewrites the current query based on the dialogue history and then utilizes the rewritten query for retrieval and response generation. RL-base rewriter utilizes RL with generation feedback to assist the training of the rewriter. ConvGQR generates potential answers for query expansion and incorporates similarity between the rewrite and gold document into the rewriter’s optimization objective.
@@
In this work, we train the query rewriter based on the pre-trained T5-base model. The dense retrieval component in the RAG system is constructed using FAISS (Johnson, Douze, and Jegou 2021) and the pre-trained embedding model msmarco-roberta-base-ance-firstp (Reimers and Gurevych 2019). The generative model is the Llama-2-13b-chat model (Touvron et al. 2023). For rewriter initialization, we employ the Adam optimizer (Kingma and Ba 2017) and train for 10k iterations with a learning rate of 5e-6 and a batch size of 16. Then we add a value head to the initialized rewriter to initialize reward models. Reward models are trained for 5 epochs using the Adam optimizer and the cosine-type scheduler. The learning rate is set to 5e-5 and the batch size is 16. During the RL phase, the learning rate is set to 1.41e-5, the batch size is 32, and the model is trained for 80k steps, with the parameter $\eta$ set to 0.05. For QReCC dataset, the hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\mu$ are set to 0.04, 0.01, 0.35, and 0.6, respectively. For TopiOCQA dataset, the values are 0.04, 0.01, 0.45, and 0.5, respectively.
@@
# Case 1:
Context (QReCC):
$q_{1}$: What are the 2 most abundant elements in the earth’s crust?
$a_{1}$: Oxygen and Silicon are the most abundant elements in the …
$q_{2}$: How much is Oxygen?
$a_{2}$: Oxygen makes up 46.6% of the Earth’s crust.
$q_{3}$: How much is Silicon?
$a_{3}$: 27.7% of the earth’s crust is silicon.
$q_{4}$: What’s the third most abundant?
$a_{4}$: Aluminum is the third most abundant element in the earth’s crust.
Current Query: What’s the atomic number?
T5-base: What’s the atomic number for silicon? (MRR: 0.0)
RL-base: What’s the atomic number for silicon? (MRR: 0.0)
MaFeRw: What’s the atomic number for aluminum? (MRR: 1.0)
Generated Responce:
T5-base: The atomic number for silicon is 14.
RL-base: The atomic number for silicon is 14.
MaFeRw: The atomic number for aluminum is 13.
Ground Truth: The atomic number of aluminum is 13.
# Case 2:
Context (QReCC):
$q_{1}$: Who won the Tour de France in 1985?
$a_{1}$: Bernard Hinault won the Tour de France in 1985.
$q_{2}$: Who won in 2019?
$a_{2}$: Egan Bernal from Team Ineos
$q_{3}$: Who was the youngest winner?
$a_{3}$: The youngest Tour de France winner was Cornet Henri,
who was 19 years old.
Current Query: Which year?
T5-base: Which year did Cornet Henri win the Tour de France in 1985?
(MRR: 0.5)
RL-base: Which year did Cornet Henri win the 1985 Tour de France?
(MRR: 0.5)
MaFeRw: Which year did Cornet Henri win the Tour de France?
(MRR: 0.5)
Generated Responce:
T5-base: Cornet Henri did not win the Tour de France in 1985.
RL-base: Cornet Henri did not win the Tour de France in 1985. The Tour
de France was first held in 1903.
MaFeRw: Cornet Henri won the Tour de France in 1904.
Ground Truth: 1904 Tour de France.
Table 4: The case study comparing with two other baselines.
@@
It can be observed that when training the rewriter using feedback from the reward model $r_{d_{+}}$ or $r_{D}$, MRR of the rewritten queries in RAG gradually increases, while ROUGE-1 shows slight decreases or fluctuations. This indicates that these two types of reward models can reliably guide the rewriter to optimize towards better retrieval performance, but are insufficient to consistently constrain the rewriter to improve generation results. When utilizing the reward model $r_{G}$ for training feedback, although there are fluctuations in ROUGE-1 and MRR metrics, there is an overall upward trend. This trend validates that the reward model can provide positive guidance to the rewriter. When directly using $m_{q}$ for feedback, ROUGE-1 exhibits a stable increase while MRR shows some variability. This indicates that the metric can, to some extent, consistently guide the rewriter towards optimizing in a manner beneficial for RAG generation. Furthermore, Figure 3 shows the changes on ROUGE-1 and MRR as iterations increase after training the rewriter with all the feedback synthetically. Both ROUGE-1 and MRR show a more stable upward trend, which further validates the stability of MaFeRw.
@@
We present a case study and a comparison with two other rewrite baselines in Table 4 to illustrate more intuitively the improvement of MaFeRw in capturing users’ needs for multi-turn dialogue RAG.

### v2

@@ Introduction
Retrieval-augmented generation (RAG) effectively addresses the issue of factual inaccuracy in large language models (LLMs) by integrating relevant retrieved information. It is widely used in various fields such as QA systems (Du and Ji 2022; Siriwardhana et al. 2023; Sharma et al. 2024), content generation (Zhao et al. 2024; Su et al. 2024; Wang et al. 2023), and virtual agents (Yang, Yue, and He 2023; Schick et al. 2023; Zhang 2023). Real-world RAG systems often rely on multi-turn dialogues, where queries include spoken ellipses and ambiguous references from dialogue contexts (Mao et al. 2023b), making it challenging for RAG systems to accurately understand user intent (Ye et al. 2024). In light of this challenge, effective query rewriting is necessary to better describe user information needs and ensure the accuracy of retrieval and generation in RAG.
@@
Currently, RAG mainly performs context-based query rewriting to improve the quality of response generation, but the resulting enhancement is minimal due to the lengthy process from query rewriting to response generation. Although context-based query rewriting can directly affect the retrieval results, this impact is difficult to transmit to the final generated results through this loose-coupling retrieval and generation framework of RAG. As shown in Figure 1, the T5 rewriter (Lin et al. 2020) trained on context-based rewrite dataset fails to effectively capture the user’s intent. To address this issue, Ma et al. try to use reinforcement learning (RL) with generation feedback to assist the rewriter. They optimize the rewriter with policy gradient and use the accuracy and ’hit’ rate of RAG inference’s answers to form the reward. However, this reward is sparse and provides little guidance in most cases, resulting in unstable training and generation results of their model. Additionally, it is inefficient to directly use the full RAG inference to obtain RL feedback. Because RL typically requires a large number of training steps to achieve good results, and each inference in RAG takes a long time. Thus, if a stable and efficient method for training the rewriter can be designed, it will undoubtedly further enhance the effectiveness of RAG.
@@
In dialogue systems, user utterances often contain omissions and ambiguous references (Fang et al. 2022; Zhang et al. 2019). Therefore, a rewriting model is needed to resolve these ambiguities in the current query and recover missing elements (e.g., anaphora) from the context to better articulate the user’s information needs. Current research on query rewriting mainly focuses on conversational search tasks, conversational QA tasks, and RAG tasks.
In conversational search tasks, researchers try to explore effective query rewriting methods to enhance the retrieval process. Some studies focus on extracting user intents from the dialogue context. For example, Lin et al. use human rewrites as labels to train a sequence-to-sequence model as the rewriter. Qian and Dou propose query rewriting explicitly highlights relevant terms in the query context, facilitating contextual modeling and thereby improving the quality of the learned contextual query embeddings. In (Ye et al. 2023), LLM is used as the rewriter and is prompted to complete a ”rewrite-then-edit” process based on the conversation history. Others incorporate retrieval feedback to enhance rewriting for retrieval tasks. Mo et al. and Mao et al. both take the feature similarity between the query and the target passage into the optimization objective when training the rewriting model and Wu et al. use the retrieval effect as a reward for reinforcement learning. The rewritten queries are encouraged to achieve better retrieval performance.
In conversational QA tasks, mainstream query rewriting studies (Kim et al. 2021; Chen et al. 2022) focus on feeding the generated results back into the rewriting model to improve the accuracy of the final answers. Kim et al. train the QA model by regularizing the consistency of generated answers, so that the answers predicted by the model according to the original questions were similar to the answers generated according to the rewritten questions. Chen et al. use the effect of generated answers as a reward for reinforcement learning, to improve the effect of rewriting on the final generated results.
In conversational RAG task, Ma et al. firstly create a pseudo-dataset for supervised warm-up training, then further optimize the rewriter using a policy gradient RL framework, using the accuracy and ’hit’ rate of LLMs’ answers to form the reward.
However, for RAG tasks, focusing solely on the retrieval process results in rewrites unsuitable for prompting LLMs during generation. Conversely, considering only the feedback from the final generated results can lead to unstable training of the rewriting model due to the sparse reward signal. Therefore, we propose MaFeRw, which combines multi-aspect reward feedbacks to produce more stable and satisfactory responses.
@@
For MaFeRw, the rewriter is trained to transform the current query $q^{k}$ into $q_{R}^{k}$ utilizing the dialogue history $h^{k}$, which enables the RAG system to generate more satisfactory results. Firstly, we use manual rewriting data to train a T5 model as the rewriter’s initialization. Secondly, we design three metrics as RL feedback: the similarity between the rewritten query $q_{R}^{k}$ and the gold document $d_{+}^{k}$, the ranking metrics of similarity between retrieved documents $D^{k}$ and the ground truth $G^{k}$, and ROUGE between the generation $g^{k}$ and the ground truth $G^{k}$. Inspired by RLAIF, we train three kinds of reward models for the above metrics on three corresponding paired data sets, as shown in Figure 2(a). Considering the obvious improvement of manual rewriting in the generation, ROUGE between manual rewrites and model rewrites is used as another feedback. Finally, we combine scores of these reward models and the rewritten ROUGE to feedback PPO training, exploring the optimal query rewriting strategy, as shown in Figure 2(b). Source code: https://github.com/yjEugenia/MaFeRw.
@@
Since user needs are embedded in the ground truth $G$, and the order of retrieved documents $D$ impacts generation quality, we hope the rewritten query can assist the retrieval model in prioritizing the document that exhibits high semantic relevance to the ground truth. Therefore, we propose a ranking metric $m_{D}$ using $G$ and $D=\{d_{1},\cdots,d_{l}\}$. Specifically, we transform the ground truth $G$ into a dense vector $v_{G}$ using the retrieval embedding model. Then the CS value is calculated between $v_{G}$ and dense vectors of retrieved documents $\{v_{d_{i}}\}_{i=1}^{l}$. For each $d_{i}$, let $1/i$ be its rank score. The metric $m_{D}$ is computed by summing the product of each document’s rank score and its CS value with $v_{G}$, formulated as $m_{D}=\Sigma_{i=1}^{l}1/i\cdot\mathrm{CS}(v_{G},v_{d_{i}})$. A higher $m_{D}$ indicates that the $d_{i}$ with greater similarity to $G$ is ranked higher, thereby facilitating the generation of better responses.
@@
The end-to-end RL approach without a reward model demands extensive training time due to the prolonged inference time of RAG and the large number of training steps required. Incorporating the complete RAG process into RL led to training durations of several days, severely limiting tuning and flexibility in later stages. To address this issue, reward models were employed to evaluate rewritten queries, effectively decoupling the evaluation from RAG process and enabling more efficient training. Inspired by RLAIF, we train three types of reward models for the aforementioned metrics on three corresponding paired datasets. The reward model is based on the T5-base model with a value head, possessing significantly fewer parameters than the generation model Llama2-13b. During RL training, the reward is derived from scores of reward models rather than using the complete RAG process, thereby reducing RAG inference time.
@@
here ”+” denotes string concatenation. When training the reward model, we adopt a method similar to Wang et al.. By following the Bradley-Terry model (Bradley and Terry 1952), we formulate a preference distribution by employing the reward model $r_{\psi}$ as outlined below:
@@
where $\eta$ is a coefficient governing the magnitude of KL penalty, the initial policy model is the initialized rewriter.
@@
We compare MaFeRw with four conversational RAG methods: raw RAG (Lewis et al. 2021), T5-base rewriter (Lin et al. 2020), RL-base rewriter (Ma et al. 2023), and ConvGQR (Mo et al. 2023). Raw RAG is the baseline model that performs retrieval and response generation without query rewriting, using the concatenation of dialogue history and the current query. T5-base rewriter uses manual rewrites to train a T5 model for query rewriting based on dialogue history. RL-base rewriter utilizes RL with generation feedback to assist the training of the rewriter. ConvGQR generates potential answers for query expansion and incorporates similarity between the rewrite and gold document into the rewriter’s optimization objective.
@@
In this work, we train the query rewriter based on the pre-trained T5-base model. The dense retrieval component in the RAG system is constructed using FAISS (Johnson, Douze, and Jegou 2021) and the pre-trained embedding model msmarco-roberta-base-ance-firstp (Reimers and Gurevych 2019). The generative model is the Llama-2-13b-chat model (Touvron et al. 2023). After initialization of the rewriter, a value head is added to the rewriter to initialize reward models. Details on hyperparameter determination are provided in the Appendix [^2] Appendix can be found at: https://arxiv.org/abs/2408.17072.
@@
Figure 4: The case study comparing with two baselines.
@@
It can be observed that when training the rewriter using feedback from the reward model $r_{d_{+}}$ or $r_{D}$, MRR of the rewritten queries in RAG gradually increases, while ROUGE-1 shows slight decreases or fluctuations. This indicates that these two types of reward models can reliably guide the rewriter to optimize towards better retrieval performance, but are insufficient to consistently constrain the rewriter to improve generation results. When utilizing the reward model $r_{G}$ for training feedback, although there are fluctuations in ROUGE-1 and MRR metrics, there is an overall upward trend. This trend validates that the reward model can provide positive guidance to the rewriter. When directly using $m_{q}$ for feedback, ROUGE-1 exhibits a stable increase while MRR shows some variability. This indicates that the metric can, to some extent, consistently guide the rewriter towards optimizing in a manner beneficial for RAG generation. Furthermore, Figure 3 shows the changes in ROUGE-1 and MRR as iterations increase after training the rewriter with synthetic feedback. Both metrics show a more stable upward trend, further validating MaFeRw’s stability.
@@
We present a case study and a comparison with two other rewrite baselines in Figure 4 to illustrate more intuitively the improvement of MaFeRw in capturing users’ needs for multi-turn dialogue RAG.
@@ Conclusion
Acknowledgments
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62406013, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-23-006) and the Fundamental Research Funds for the Central Universities.
Appendix A Details of Experimental Settings
The experiments are conducted on one Nvidia A100 40G GPU. We implement all models by PyTorch (Paszke et al. 2019), LangChain (Chase 2022) and Huggingface’s Transformers (Wolf et al. 2019). For rewriter initialization, we employ the Adam optimizer (Kingma and Ba 2017) and train for 10k iterations with a learning rate of 5e-6 and a batch size of 16. Then we add a value head to the initialized rewriter to initialize reward models. Reward models are trained for 5 epochs using the Adam optimizer and the cosine-type scheduler. The learning rate is set to 5e-5 and the batch size is 16. During the RL phase, the learning rate is set to 1.41e-5, the batch size is 32, and the model is trained for 80k steps, with the parameter $\eta$ set to 0.05. For QReCC dataset, the hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\mu$ are set to 0.04, 0.01, 0.35, and 0.6. For TopiOCQA dataset, the values are 0.04, 0.01, 0.45, and 0.5, respectively.
For the selection of hyperparameters in the reward function, we first refer to the improvement of retrieval and generation performance when a certain metric is used alone, and make the initial assignment according to the proportion, then use grid search to fine-tune and select the optimal parameters based on experimental results. To better compare with other baselines like ConvGQR, the maximum generation length is set to 32. QReCC focuses on the query rewriting problem within conversational scenarios by approaching the human-rewritten query. Thus, it provides an oracle query for each conversation turn. TopiOCQA focuses on the challenge of the topic switch under conversational settings, whose sessions are longer than QReCC and thus present more difficulties for query reformulation. Different from QReCC, it does not provide human-rewritten queries. The publisher of TopiOCQA finetunes the T5 model on a rewrite of QReCC and uses it to generate a rewrite of TopiOCQA. Since this is obtained based on manual rewriting training, we take it as the label of TopiOCQA representing manual rewriting. During the retrieval process, the number of retrieved articles was set to 5.
When training the reward model, we randomly selected 48% of the whole training set as the reward model training set and 12% as the test set instead of using all the data for the training of the reward model. This can better verify whether the reward model is overfitting. Due to hardware limitations, we filtered the document collections of QReCC and TopiOCQA datasets, retaining only the positive documents from the training and test sets, along with randomly sampled negative documents, while also removing any corrupted files.
We implement baselines based on our experimental setting and their open-source code and material. For the normal evaluation, we train T5-base, RL-base, and ConvGQR on the corresponding datasets rather than using external resources.